OpenStudy (anonymous):

What is the equation of the line that passes through (–2, 3) and is parallel to 2x + 3y = 6?

OpenStudy (anonymous):

To answer that question, first, make the 2x + 3y = 6 in slope form.

OpenStudy (anonymous):

so y=2x+6?

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

sorry im terrible at math

OpenStudy (anonymous):

it's okay. :) to make that in slope form... 2x + 3y = 6 -- it will be like this.. 3y = -2x + 6 y = -2/3 x + 2 got it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Good. :) So, what's the slope of the equation?

OpenStudy (anonymous):

-2/3

OpenStudy (anonymous):

Correct. :) Then use this formula to find the equation of the line which is parallel to 2x + 3y = 6